OpenStudy (anonymous):

7th grade math help plzzz

OpenStudy (anonymous):

Which expression is equivalent to this one? –7(5 + 10) A. (–7) · 5 + (–7) · 10 B. (–7) · 5 + 10 C. (–7) + 5 + 10 D. (–7) · 5 · 10

OpenStudy (anonymous):

Its A

OpenStudy (anonymous):

Rewrite the expression by using grouping symbols. y • 4 + y • z A. y (4 + z) B. y • (4 + y) • z C. z(y • 4) D. 4(y + z)

OpenStudy (anonymous):

i see a lot of ?????????????

OpenStudy (anonymous):

leave then come back in you could see it better:(

OpenStudy (anonymous):

ok much better

OpenStudy (anonymous):

its A again XD

OpenStudy (anonymous):

What is the solution to the equation? 4(3 + 7) = 4 · x + 4 · 7 A. x = 3 B. x = 4 C. x = 7 D. x = 10

OpenStudy (anonymous):

x=3 A again, they should have put these answers in a different space XD

OpenStudy (anonymous):

9 * 72 = 9 * ____________+ 9 * 2 =____________

OpenStudy (anonymous):

27 for the first blank, and 648 for the second

OpenStudy (anonymous):

I got a 60

OpenStudy (anonymous):

for the first blank?? i thought i was kinda off

OpenStudy (anonymous):

no no, its 35 for sure ok